Mason Office Services color copies for various customers. The company has gathered the following data:

Month                         # Copies                      Copy Department Costs

July                             17,000                                     $1,445

August                        21,250                                       1,685

September                   22,950                                       1,785

October                       18,700                                       1,530

November                   20,400                                       1,615

December                   25,500                                       2,040

Required:

(a) Enter the above data onto an Excel worksheet. Then, using the High-Low method, build the applicable regression equation, and

                        (i)    Solve for “f.”

                        (ii)   Solve for “v”

                        (iii) Estimate Copy Department Costs, assuming the number of copies is 23,000.

(b) On the same worksheet, using Excel’s Least Squares Regression application,

                        (i) Solve for “f.”

                        (ii) Solve for “v.”

                        (iii) Estimate Copy Department Costs, assuming the number of copies is 23,000.

                        (iv) Based on the r2 measure (coefficient of determination), explain what it indicates about how accurate (or inaccurate) the Least Squares equation is at predicting total costs.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Mason Office Services
Answer a
High Low method
Variable cost= (Highest activity cost- Lowest activity cost)/ (Highest activity units- Lowest activity units)
Month Costs Copies
Highest activity December                  2,040.00                   25,500.00
Lowest activity July                  1,445.00                   17,000.00
Difference                      595.00                     8,500.00
Variable cost per unit                          0.07 This is answer ii.
Fixed cost Highest activity cost- (Highest units * Variable cost per unit)
Fixed cost 2040-(25500*0.07)
Fixed cost                  255.00 This is answer i.
23000 copies
Copies             23,000.00
Variable cost per unit                       0.07
Total Variable cost               1,610.00
Fixed cost                  255.00
Estimated total cost               1,865.00 This is answer iii.
Answer b
Least Squares Regression
The equation of a line is y=mx+b
X= number of copies
y=total cost
m= (N Σ(xy) − Σx Σy)/ (N Σ(x2) − (Σx)2)
Month Copies (x) Costs (y) x2 xy
July             17,000.00                  1,445.00         289,000,000.00        24,565,000.00
August             21,250.00                  1,685.00         451,562,500.00        35,806,250.00
September             22,950.00                  1,785.00         526,702,500.00        40,965,750.00
October             18,700.00                  1,530.00         349,690,000.00        28,611,000.00
November             20,400.00                  1,615.00         416,160,000.00        32,946,000.00
December             25,500.00                  2,040.00         650,250,000.00        52,020,000.00
Total          125,800.00                10,100.00     2,683,365,000.00      214,914,000.00
N= number of months= 6.
m= (N Σ(xy) − Σx Σy)
(N Σ(x2) − (Σx)2)
m= 6*214914000- 125800*10100
6*2683365000- 125800^2
m= 1289484000- 1270580000
16100190000-15825640000
m=     18,904,000.00
274,550,000.00
m=                       0.07
b= (Σy − m Σx)/N
b= 10100- 0.07*125800
                      6.00
b= 10100- 8661.89473684211
                      6.00
b=               1,438.11
                      6.00
b=                  239.68
The equation of a line is y=mx+b
y= 0.07x+239.68
Answer i f= $ 239.68
Answer ii v= $ 0.07
Answer iii
Cost of 23000 copies 0.07*23000+239.68
Cost of 23000 copies               1,823.34